# OriStitch: A Machine Embroidery Workflow to Turn Existing Fabrics into Self-Folding 3D Textiles

### v1

@@ 6. Applications
We demonstrate OriStitch along a series of different applications. In fabricating these, we note that the boundary edge of the model is not as tightly folded as the other hinges, this is a common problem for folded geometry as identified by Panetta et al. (2021), to this end we sew a 3D printed rim of the geometry along the boundary of the mesh to provide outer faces with the required supporting force.
6.1. Cap: custom fashion
Fabric pleats have long been used in fashion items like cloth. Inspired by the tradition, we designed a cap. The design consists of 211 faces and 303 hinges. It was fabricated with 22 stages of hooping, the result is shown in Figure LABEL:fig:cap-closeup.
6.2. Vase Cover: shape-matching existing geometry
Textiles are often used in interior decoration creating visual and textual effects as well as other functions like protection. The textile cover is designed to fit the organic ceramic vase making it visually appealing while providing protection. The vase cover consists of 100 faces and 140 hinges, fabricating with 11 hoops.
6.3. Hand Bag: enclosing volumes
The choice of fashion accessories reflects personal style and preferences, enabling individuals to express their unique tastes and create a distinctive look. This custom-shaped bag features 232 faces and 338 hinges, fabricated using 17 hoops.
@@ 7.1. Forces at the Boundary
Our conversion tool assumes the faces of the model are rigid, this is a fair assumption as there are many stitches and threads spanning each face making them relatively stiff. However, the faces and edges at the boundary of the sheet are still as flexible as the underlying fabric, we therefore reinforce the application models with a 3D printed rim. It would be interesting to add an additional stitch along the boundary to reinforce, or even stitch a channel to insert a non-fabric material (metal or plastic filament) in the rim after actuation. This will support the other (non-boundary) faces as well.
@@ 7.2. Actuation Method
In finding the ideal actuation method, we have tried many options ranging from hair straighteners to ovens to oil baths. We generally found that OriStitch hinges are best actuated in boiling water simultaneous with heat gun post processing in sequence. The method ensures effective actuation with minimal manual effort. It is a typical problem in Origami that while the crease pattern will uniquely fold into the target shape, it is not always known that valid folding motions exist, let alone that they can all be folded simultaneously. Interesting follow-up work could investigate this path and/or explore other actuation methods like a heat gun on a robotic arm similar to the set up in FormFab (Mueller et al., 2019).
@@ 7.3. Need of Guidance for Multi-hooping
While OriStitch automatically generates the stitch and cutting files, there is still manual labor operating embroidery machines. Unless users have an industrial-type of machine, they will likely revert to multi-hooping, which requires elaborate alignment effort and frequent rewiring of threads. Right now, fabricating an OriStitch pattern requires intimate knowledge of the machine and materials. More work should be done to reduce this effort for machine-embroidery in general, most steps are repetitive, but easy to mix up and it requires experience to oversee the machine to see when something is off. Better computer vision tools as well as process guidance could go a long way to make this process more accessible.
@@ 7.4. Size of Stitch Patterns
Multi-hooping proved to be instrumental to make larger sheets because of the limitations of the size of embroidery machines. However, especially for models with a lot of detail, the stitch patterns get very large. Some models from our technical evaluation exceed the size of the laser cutter as well. While it is possible to lasercut and stitch multiple sheets together, we note that depending on the used fabric, there is a limit to how much each hinge can carry. A future version of OriStitch should use the material properties and hinge layout to simulate how large a structure it supports, and maybe even locally reinforce similar to FastForce (Abdullah et al., 2021) for lasercutting.
@@ 8. Conclusion
We presented OriStitch, a machine-embroidery workflow, and software tool to create 3D self-folding textiles. We introduce a specific hinge design, which self-folds, is flexible in width, and fabricates reliably in a range of different materials. We have shown how this method allows us to create the cap of Figure 1. We then evaluated our conversion and found that 23/28 models found in related papers are supported by our process, and identify how to extend the tool to address issues in the remaining five models.
The OriStitch workflow makes creating intricate 3D textiles more accessible. In the future, we hope to see more workflows to reduce the labor required for fabrication of expressive textile objects. Such custom fabrication used to be prohibitively expensive and solely accessible in high-fashion. Accessible workflows encourage exploration and inclusive design for both high-end and hobbyist users.

### v2

@@ 6. Applications
We demonstrate OriStitch through a series of applications. In fabricating these, we note that the boundary edge of the model is not as tightly folded as the other hinges, this is a common problem for folded geometry as identified by Panetta et al. (2021), to this end we sew a 3D printed rim of the geometry along the boundary of the mesh to provide outer faces with the required supporting force. Figure 17 shows the different application model, (a-c) their conversion from 3D to stitch pattern and (d) the final fabricated artifacts.
Cap: custom fashion. Fabric pleats have long been used in fashion items like clothes. Inspired by the tradition, we designed a cap. The design consists of 211 faces and 303 hinges. It was fabricated with 11 stages of hooping. This is the cap featured in Figure one.
Vase Cover: shape-matching existing geometry. Textiles are often used in interior decoration, creating visual and textual effects as well as other functions like protection. The textile cover is designed to fit the organic ceramic vase, making it visually appealing while providing protection. The vase cover consists of 100 faces and 140 hinges, fabricated with 5 hoops. This model further demonstrates another advantage of the additive nature of Oristitch: that is combines excellently with other textile fabrication. Here the dots have been embroidered on the material beforehand.
Hand Bag: enclosing volumes. The choice of fashion accessories reflects personal style and preferences, enabling individuals to express their unique tastes and create a distinctive look. This custom-shaped bag features 232 faces and 338 hinges, fabricated using 12 hoops.
@@ 7.1. Forces at the Boundary
Our conversion tool assumes the faces of the model are rigid; this is a fair assumption as there are many stitches and threads spanning each face, making them relatively stiff. However, the faces and edges at the boundary of the sheet are still as flexible as the underlying fabric; we therefore reinforce the application models with a 3D printed rim. It would be interesting to add an additional stitch along the boundary to reinforce, or even stitch a channel to insert a non-fabric material (metal or plastic filament) in the rim after actuation. This will support the other (non-boundary) faces as well.
7.2. Active thread
For the purpose of accessibility of our approach, we went with an active thread that is available on the shelf and shrinks by 30% and generates limited forces. Developing shrinking threads with higher shrinkage and stronger force would simplify the hinge designs, speed up fabrication process, and make it applicable to heavier materials. Additionally, the thread we are using is one-time shrinking; using revertible shrinking threads would open up additional possibilities. Researchers have explored the fabrication of custom fibers (Forman et al., 2023), but this exceeds the scope of this paper.
@@ 7.3. Actuation Method
In finding the ideal actuation method, we have tried many options, ranging from hair straighteners to ovens to oil baths. We generally found that OriStitch hinges are best actuated in boiling water simultaneous with heat gun post processing in sequence. The method ensures effective actuation with minimal manual effort. It is a typical problem in Origami that while the crease pattern will uniquely fold into the target shape, it is not always known that valid folding motions exist, let alone that they can all be folded simultaneously. Interesting follow-up work could investigate this path and/or explore other actuation methods like a heat gun on a robotic arm similar to the set up in FormFab (Mueller et al., 2019).
@@ 7.4. Need of Guidance for Multi-hooping
While OriStitch automatically generates the stitch and cutting files, there is still manual labor operating embroidery machines. Unless users have an industrial-type of machine, they will likely revert to multi-hooping, which requires elaborate alignment effort and frequent rewiring of threads. Right now, fabricating an OriStitch pattern requires intimate knowledge of the machine and materials. More work should be done to reduce this effort for machine-embroidery in general; most steps are repetitive, but easy to mix up, and it requires experience to oversee the machine to see when something is off. Better computer vision tools, as well as process guidance, could go a long way to make this process more accessible.
@@ 7.5. Size of Stitch Patterns
Multi-hooping proved to be instrumental in making larger sheets because of the limitations of the size of embroidery machines. However, especially for models with a lot of detail, the stitch patterns get very large. Some models from our technical evaluation exceed the size of the laser cutter as well. While it is possible to laser-cut and stitch multiple sheets together, we note that depending on the fabric used, there is a limit to how much each hinge can carry. A future version of OriStitch should use the material properties and hinge layout to simulate how large a structure it supports, and maybe even locally reinforce, similar to FastForce (Abdullah et al., 2021) for lasercutting.
7.6. Advancing Embroidery Machines
In this paper, we aimed to fabricate all models using an accessible home embroidery machine from Brother. We opted for this to minimize dependency on some rare machine properties and thus reduce the replicability of our approach. Obviously, larger machines will reduce the required number of hoops and time for fabrication; furthermore, machines with multiple needles allow fabricating more layers at once, and machines equipped with computer vision techniques further simplify the registration process. We did experiment with more advanced machines to confirm those capabilities. An extension of machines that would further increase the capabilities of our method would be integration with a laser cutter. Many machines have the ability to be equipped with cutting needles, but the resolution and orientation of the blade dramatically reduce its value for precision embroidery.
@@ 8. Conclusion
We presented OriStitch, a machine-embroidery workflow and software tool to create 3D self-folding textiles. We introduce a specific hinge design, which self-folds, is flexible in width, and fabricates reliably in a range of different materials. We have shown how this method allows us to create the cap of Figure 1. We evaluated our conversion and found that 26/28 models found in related papers are supported by our process, and identified how to extend the tool to address issues in the remaining five models.
The OriStitch workflow makes creating intricate 3D textiles more accessible. In the future, we hope to see more workflows to reduce the labor required for the fabrication of expressive textile objects. Such custom fabrication used to be prohibitively expensive and solely accessible in high-fashion. Accessible workflows encourage exploration and inclusive design for both high-end and hobbyist users.